40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6910     Entry time: Tue Jul 3 20:51:06 2012
Author: yuta 
Type: Update 
Category: IOO 
Subject: MC in vacuum is back 

MC came back to the state as it was before the vent.

What I did:
  1. Removed beam attenuating setup on PSL table(see elog #6892).

  2. Removed 100% reflection mirror before the MC reflection PD and put 10% BS back, so that we can have MC WFS. Also, I changed C1:IOO-MC_RFPD_DCMON.HOPR to 5.

  3. Removed autolockMCmain40m_low_power from crontab on op340m, and put autolockMCmain40m again.

  4. Aligned MC and ran /opt/rtcds/caltech/c1/scripts/MC/WFS/WFS_FilterBank_offsets to adjust WFS offsets.

  5. Measured beam spot positions. They looked same as before the vent.

# filename    MC1pit    MC2pit    MC3pit    MC1yaw    MC2yaw    MC3yaw    (spot positions in mm)
./dataMCdecenter/MCdecenter201206290135.dat    2.914584    4.240889    2.149244    -7.117336    -1.494540    4.955329    before vent
./dataMCdecenter/MCdecenter201207011253.dat    3.294659    3.416584    2.620511    -6.691800    -3.164084    4.806517    after vent
./dataMCdecenter/MCdecenter201207032009.dat    3.737099    3.994597    3.087857    -6.442053    -0.992543    4.714607    after pumping (now)

  6. I also turned on high voltage power supplies for input and output PZTs

  7. Below is captured Sensoray images of the current state.
ALL_1025408289.bmp


Next:
  I will go on to check if IFO works as it was before or not, but I think we should center MC beam spot positions and see if we can avoid clipping in the near future.

ELOG V3.1.3-